Hi, not sure if this question is in the correct area but please can some one answer it?

A ball is dropped from a building. The distance, D metres, that it falls in t seconds is given by
D=5t squared
a. Find the distance fallen in 3 seconds
B. Find the time to fall 100 metres

Can I please have all the working out, it helps me understand how to complete the question.
This is much appreciated

Answer by Alan3354(69443)

A ball is dropped from a building. The distance, D metres, that it falls in t seconds is given by
D=5t^2
a. Find the distance fallen in 3 seconds
Sub 3 for t
D = 5*3^2 = 45 meters
---------------
B. Find the time to fall 100 metres
100 = 5t^2
t^2 = 20
t =~ 4.472 seconds